Home‎ > ‎Crew 133‎ > ‎

116-GreenHabReport

XO Gartrelle to Cap Comm
The green hab is operating normally

The fan was set to 60 degrees at 8AM
The fan is operating normally at 68 degrees at 5PM

The heater was turned down to 40 degrees F at 8AM
The heater was turned up to 75 degrees F at 5PM

The pump is operating normally.

The herb seedlings still have not sprouted. The Mars shoot is still looking
healthy.
The herb seedlings and the Mars shoot received water this morning.

The hydroponic lettuce sprouts received no water today. Eleven of the
twelve shoots look healthy at 5PM.  The droopy mint lettuce shoot from
yesterday seems to have recovered. It is the upper one in the photo of the
single hydroponic bag. At 5PM the hydroponic soil appeared damp. The
sprouts will be watered tomorrow at 9AM.

They remaining lettuce seedlings look very healthy. The pepper seedlings
have not sprouted as yet. All the seedlings in root cubes received water at
8AM.
All the seedlings were returned to the upper hab at 5PM.

No other plants received water today.

Greenhab readings for 5PM today are as follows:

                    Current         Min          Max
Temperature        82 Degrees F    48 Degrees F 100 Degrees F
Humidity           14%             19 %          10%

Meter            0.94 KwH
Duration        22.55 Hours
Overflow         0 Liters    Removed Today - Checked at 8AM and at 5PM


Photos from today are attached.

Regards,
End Transmission-